# Graphon Quantum Filtering systems

### v1

@@ 4 Mean-field graphon quantum filtering systems
    $\Big{\|}\tilde{\gamma}_{t,u}-\hat{\gamma}_{t,u}\Big{\|}^{2}_{2}=\Big{\|}\int_{0}^{t}\Big{(}-\mathrm{i}\big{[}\tilde{H},\tilde{\gamma}_{s,u}-\hat{\gamma}_{s,u}\big{]}-\mathrm{i}\big{[}\int_{0}^{1}\tilde{w}(u,v)A^{\mathbb{E}[\tilde{\gamma}_{s,v}]}\mathrm{d}v,\tilde{\gamma}_{s,v}\big{]}$
    $+\mathrm{i}\big{[}\int_{0}^{1}\hat{w}(u,v)A^{\mathbb{E}[\hat{\gamma}_{s,v}]}\mathrm{d}v,\hat{\gamma}_{s,u}\big{]}\Big{)}\mathrm{d}s+\int_{0}^{t}\Big{(}L(\tilde{\gamma}_{s,u}-\hat{\gamma}_{s,u})L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,(\tilde{\gamma}_{s,u}-\hat{\gamma}_{s,u})\}\Big{)}\mathrm{d}s$
    $+\sqrt{\eta}\int_{0}^{t}\Big{(}L(\tilde{\gamma}_{s,u}-\hat{\gamma}_{s,u})+(\tilde{\gamma}_{s,u}-\hat{\gamma}_{s,u})L^{\dagger}-\mathrm{t}\mathrm{r}(K\tilde{\gamma}_{s,u})\tilde{\gamma}_{s,u}+\mathrm{t}\mathrm{r}(K\hat{\gamma}_{s,u})\hat{\gamma}_{s,u}\Big{)}\mathrm{d}W_{s}^{u}\Big{\|}_{2}^{2}$
    $\leq$ $2\Big{\|}\int_{0}^{t}\Big{(}\big{[}\tilde{H},\tilde{\gamma}_{s,u}-\hat{\gamma}_{s,u}\big{]}\Big{)}\mathrm{d}s\Big{\|}_{2}^{2}+2\Big{\|}\int_{0}^{t}\Big{(}\big{[}\int_{0}^{1}\tilde{w}(u,v)A^{\mathbb{E}[\tilde{\gamma}_{s,v}-\hat{\gamma}_{s,v}]}\mathrm{d}v,\tilde{\gamma}_{s,v}\big{]}\Big{)}\mathrm{d}s\Big{\|}_{2}^{2}$
    $+2\Big{\|}\int_{0}^{t}\Big{(}\big{[}\int_{0}^{1}\hat{w}(u,v)A^{\mathbb{E}[\hat{\gamma}_{s,v}]}\mathrm{d}v,(\hat{\gamma}_{s,u}-\tilde{\gamma}_{s,u})\big{]}\Big{)}\mathrm{d}s\Big{\|}_{2}^{2}$
    $+2\Big{\|}\int_{0}^{t}\Big{(}\big{[}\int_{0}^{1}(\tilde{w}(u,v)-\hat{w}(u,v))A^{\mathbb{E}[\hat{\gamma}_{s,v}]}\mathrm{d}v,\tilde{\gamma}_{s,u}\big{]}\Big{)}\mathrm{d}s\Big{\|}_{2}^{2}$
@@
    $\mathbb{E}\Big{[}\Big{\|}\tilde{\gamma}_{t,u}-\hat{\gamma}_{t,u}\Big{\|}^{2}_{2}\Big{]}$
    $\leq C\mathbb{E}\Big{[}\int_{0}^{t}\Big{\|}\tilde{\gamma}_{s,u}-\hat{\gamma}_{s,u}\Big{\|}_{2}^{2}\mathrm{d}s\Big{]}+\mathbb{E}\Big{[}\Big{\|}\int_{0}^{t}\Big{(}\big{[}\int_{0}^{1}(\tilde{w}(u,v)-\hat{w}(u,v))A^{\mathbb{E}[\hat{\gamma}_{s,v}]}\mathrm{d}{v},\tilde{\gamma}_{s,u}\big{]}\Big{)}\mathrm{d}s\Big{\|}_{2}^{2}\Big{]},$
    $\leq C\Big{(}\int_{0}^{t}\mathbb{E}\Big{[}\Big{\|}\tilde{\gamma}_{s,u}-\hat{\gamma}_{s,u}\Big{\|}_{2}^{2}\Big{]}\mathrm{d}s+\int_{0}^{t}\int_{0}^{1}\Big{\|}(\tilde{w}(u,v)-\hat{w}(u,v))A^{\mathbb{E}[\hat{\gamma}_{s,v}]}\Big{\|}_{2}^{2}\mathrm{d}v\mathrm{d}s\Big{)},$
@@
    $\leq C\Big{(}\int_{0}^{t}\mathbb{E}\Big{[}\Big{\|}\tilde{\gamma}_{s,u}-\hat{\gamma}_{s,u}\Big{\|}_{2}^{2}\Big{]}\mathrm{d}s+T\big{\|}a\big{\|}_{2}^{2}\int_{0}^{1}\Big{|}(\tilde{w}(u,v)-\hat{w}(u,v))\Big{|}^{2}\mathrm{d}v\Big{)}.$
@@
    $\mathbb{E}\Big{[}\int_{0}^{1}\Big{\|}\tilde{\gamma}_{t,u}-\hat{\gamma}_{t,u}\Big{\|}^{2}_{2}\mathrm{d}u\Big{]}\leq$ $C\Big{(}\int_{0}^{t}\mathbb{E}\Big{[}\int_{0}^{1}\Big{\|}\tilde{\gamma}_{s,u}-\hat{\gamma}_{s,u}\Big{\|}_{2}^{2}\mathrm{d}u\Big{]}\mathrm{d}s$
@@
Remark 3 (Strong Solution).
In the case of perfect measurement where detector efficiency $\eta=1$, homogeneous graphon $w\equiv 1$ and bounded coupling operator $L$, the well-posedness in the strong sense was proved by a different method in [44]. An extension to the unbounded coupling operator was recently provided in [45]. For others results concerning well-posedness of stochastic Schrödinger equation one can see [54, 33, 36] in infinite dimensional setting and [57] for density operator equation in finite dimensional setting.
@@ 5 Propagation of chaos: Block-wise case
For clarity, we slightly re-adjust the framework and notation. Specifically, we consider $c$-packets of $N$ quantum particles interacting through a graph $G_{cN}=(V_{cN},E_{cN})$.
@@
$\mathrm{d}\boldsymbol{\rho}^{cN}_{t}$ $=-\mathrm{i}\Big{[}\sum_{i_{l}=1_{1}}^{c_{N}}\mathbf{\tilde{H}}_{i_{l}}\!+\frac{1}{2cN}\sum_{i_{l}=1_{1}}^{c_{N}}\sum_{j_{l^{\prime}}=1_{1}}^{c_{N}}\xi_{j_{l^{\prime}}i_{l}}\mathbf{A}_{j_{l^{\prime}}i_{l}},\boldsymbol{\rho}^{cN}_{t}\Big{]}\mathrm{d}t+\sum_{i_{l}=1_{1}}^{c_{N}}\big{(}\mathbf{L}_{i_{l}}\boldsymbol{\rho}^{cN}_{t}\mathbf{L}_{i_{l}}^{\dagger}-\frac{1}{2}\{\mathbf{L}_{i_{l}}^{\dagger}\mathbf{L}_{i_{l}},\boldsymbol{\rho}^{cN}_{t}\}\big{)}\mathrm{d}t$
$\qquad+\sum_{i_{l}=1_{1}}^{c_{N}}\Big{(}\mathbf{L}_{i_{l}}\boldsymbol{\rho}^{cN}_{t}+\boldsymbol{\rho}^{cN}_{t}\mathbf{L}_{i_{l}}^{\dagger}-\mathrm{t}\mathrm{r}\big{(}(\mathbf{L}_{i_{l}}+\mathbf{L}_{i_{l}}^{\dagger})\boldsymbol{\rho}^{cN}_{t}\big{)}\boldsymbol{\rho}^{cN}_{t}\Big{)}\mathrm{d}W_{t}^{i_{l}}.$ (5.1)
@@
$\mathbf{O}_{i_{l}}$ $:=\overbrace{\mathbf{1}\otimes\dots\otimes\mathbf{1}}^{N(i-1)\text{ times}}\otimes\overbrace{\mathbf{1}\otimes\dots\otimes\underbrace{O}_{l-\text{th}}\otimes\dots\otimes\mathbf{1}}^{N\text{ times}}\otimes\overbrace{\mathbf{1}\otimes\dots\otimes\mathbf{1}}^{(c-i+1)N\text{ times}},$
@@
The interaction strength coefficient $\xi_{j_{l^{\prime}}i_{l}}:=\xi_{j_{l^{\prime}}i_{l}}^{cN}$ is derived from a step kernel function ${w}^{G_{cN}}$, and is constructed in two following ways:
@@
$w(u,v)$ :=$:=\begin{cases}w_{ji},\quad(u,v)\in I_{j}\times I_{i},\\ 0,\quad\quad\,\,\,\textrm{otherwise.}\end{cases}$
@@
If the initial state is pure, i.e., $\boldsymbol{\rho}^{cN}_{0}=\ket{\boldsymbol{\Psi}_{0}^{cN}}\bra{\boldsymbol{\Psi}_{0}^{cN}}$ |, then, by Proposition 1, then the state remains pure i.e $\boldsymbol{\rho}^{cN}_{t}=\ket{\boldsymbol{\Psi}_{t}^{cN}}\bra{\boldsymbol{\Psi}_{t}^{cN}}$ | and can be written in the following pure state representation:
@@
Remark 4.
The case where the blocks are not of equal size can be treated similarly. Indeed, we can consider $c$ packets, each containing $N_{i}$ particles, for $i\in[c]$, such that $\sum_{i=1}^{c}N_{i}=N,$ and $\lim_{N\to\infty}\frac{N_{i}}{N}=O(1)$. For simplicity of notation and exposition, we focus here on the case of equally-sized blocks, i.e., $N_{i}=N$ for all $i$.
@@
Proposition 2 (Class‐wise symmetry).
@@
$\textrm{Law}\Big{(}\boldsymbol{\Psi}_{0}^{cN}\bigl{(}x_{1_{1}},\dots,x_{c_{N}}\bigr{)}\Big{)}$ $=\textrm{Law}\Big{(}{\Psi}_{0}^{cN}\Bigl{(}x_{1_{\sigma^{1}(1)}},\dots,x_{1_{\sigma^{1}(N)}},\dots,x_{c_{\sigma^{c}(1)}},\dots,x_{c_{\sigma^{c}(N)}}\Bigr{)}\Big{)},$ (5.2)
for all permutations $\sigma^{1},\dots,\sigma^{c}\in\mathfrak{S}_{N}$. Then this symmetry is preserved in distribution by the $cN$-body Belavkin Equation (5.1) if the interaction coefficients $\{\xi_{j_{l^{\prime}}i_{l}}\}$ are invariants in law:
@@
Under Assumption 1 and Assumption 2, (5.1) describes the dynamics of $c$-class of Bosons.
@@ 5.1 Limiting system
$\mathrm{d}\gamma_{t,u}$ $=-\mathrm{i}\Big{[}\tilde{H}+\int_{0}^{1}w(u,v)A^{\mathbb{E}[\gamma_{t,v}]}\mathrm{d}v,\gamma_{t,u}\Big{]}\mathrm{d}t+\Big{(}L\gamma_{t,u}L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\gamma_{t,u}\}\Big{)}\mathrm{d}t$
@@
$\int_{0}^{1}{w}\Big{(}\frac{j-1}{c},v\Big{)}A_{j}^{\mathbb{E}[\boldsymbol{\gamma_{t,v}}]}\mathrm{d}v$ $=\sum_{i=1}^{c}\int_{0}^{1}\mathbb{I}_{\Big{\{}\frac{i-1}{c}\leq v<\frac{i}{c}\Big{\}}}{w}\Big{(}\frac{j-1}{c},v\Big{)}A_{j}^{\mathbb{E}[\boldsymbol{\gamma_{t,v}}]}\mathrm{d}v,$
$=\sum_{i=1}^{c}\int_{0}^{1}\mathbb{I}_{\Big{\{}\frac{i-1}{c}\leq v<\frac{i}{c}\Big{\}}}{w}\Big{(}\frac{j-1}{c},\frac{i-1}{c}\Big{)}A_{j}^{\mathbb{E}[\boldsymbol{\gamma_{t,v}}]}\mathrm{d}v,$
$=\sum_{i=1}^{c}\int_{\frac{i-1}{c}}^{\frac{i}{c}}{w}\Big{(}\frac{j-1}{c},\frac{i-1}{c}\Big{)}A_{j}^{\mathbb{E}[\boldsymbol{\gamma_{t,i}}]}\mathrm{d}v,$
$=\frac{1}{c}\sum_{i=1}^{c}{w}\Big{(}\frac{j-1}{c},\frac{i-1}{c}\Big{)}A_{j}^{\mathbb{E}[\boldsymbol{\gamma_{t,i}}]},$
$=\frac{1}{c}\sum_{i=1}^{c}w_{ji}A_{j}^{\mathbb{E}[\boldsymbol{\gamma_{t,i}}]}.$
@@
$\mathrm{d}\gamma_{t,j}$ $=-\mathrm{i}\Big{[}\tilde{H}+\frac{1}{c}\sum_{i=1}^{c}w_{ji}A^{\mathbb{E}[\gamma_{t,i}]},\gamma_{t,j}\Big{]}\mathrm{d}t+\big{(}L\gamma_{t,j}L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\gamma_{t,j}\}\big{)}\mathrm{d}t$
$\qquad\qquad+\Big{(}L\gamma_{t,j}+\gamma_{t,j}L^{\dagger}-\mathrm{t}\mathrm{r}\big{(}(L+L^{\dagger})\gamma_{t,j})\gamma_{t,j}\big{)}\Big{)}\mathrm{d}W_{t}^{j},\;\;j\in[c].$ (5.3)

### v3

@@ 4 Mean-field graphon quantum filtering systems
    $\Big\|\tilde{\gamma}_{t,u}-\hat{\gamma}_{t,u}\Big\|^{2}_{2}=\Big\|\int_{0}^{t}\Big(-\mathrm{i}\big[\tilde{H},\tilde{\gamma}_{s,u}-\hat{\gamma}_{s,u}\big]-\mathrm{i}\big[\int_{0}^{1}\tilde{w}(u,v)A^{\mathbb{E}_{\mathbb{P}}[\tilde{\gamma}_{s,v}]}\mathrm{d}v,\tilde{\gamma}_{s,v}\big]$
    $\qquad\qquad\qquad+\mathrm{i}\big[\int_{0}^{1}\hat{w}(u,v)A^{\mathbb{E}_{\mathbb{P}}[\hat{\gamma}_{s,v}]}\mathrm{d}v,\hat{\gamma}_{s,u}\big]\Big)\mathrm{d}s$
    $\qquad+\int_{0}^{t}\Big(L(\tilde{\gamma}_{s,u}-\hat{\gamma}_{s,u})L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,(\tilde{\gamma}_{s,u}-\hat{\gamma}_{s,u})\}\Big)\mathrm{d}s$
    $\qquad+\sqrt{\eta}\int_{0}^{t}\Big(L(\tilde{\gamma}_{s,u}-\hat{\gamma}_{s,u})+(\tilde{\gamma}_{s,u}-\hat{\gamma}_{s,u})L^{\dagger}-\mathrm{t}\mathrm{r}(K\tilde{\gamma}_{s,u})\tilde{\gamma}_{s,u}+\mathrm{t}\mathrm{r}(K\hat{\gamma}_{s,u})\hat{\gamma}_{s,u}\Big)\mathrm{d}W_{s}^{u}\Big\|_{2}^{2}$
    $\leq\;$ $2\Big\|\int_{0}^{t}\big[\tilde{H},\tilde{\gamma}_{s,u}-\hat{\gamma}_{s,u}\big]\mathrm{d}s\Big\|_{2}^{2}+2\Big\|\int_{0}^{t}\Big[\int_{0}^{1}\tilde{w}(u,v)A^{\mathbb{E}_{\mathbb{P}}[\tilde{\gamma}_{s,v}-\hat{\gamma}_{s,v}]}\mathrm{d}v,\tilde{\gamma}_{s,v}\Big]\mathrm{d}s\Big\|_{2}^{2}$
    $+2\Big\|\int_{0}^{t}\Big[\int_{0}^{1}\hat{w}(u,v)A^{\mathbb{E}_{\mathbb{P}}[\hat{\gamma}_{s,v}]}\mathrm{d}v,(\hat{\gamma}_{s,u}-\tilde{\gamma}_{s,u})\Big]\mathrm{d}s\Big\|_{2}^{2}$
    $+2\Big\|\int_{0}^{t}\Big[\int_{0}^{1}(\tilde{w}(u,v)-\hat{w}(u,v))A^{\mathbb{E}_{\mathbb{P}}[\hat{\gamma}_{s,v}]}\mathrm{d}v,\tilde{\gamma}_{s,u}\Big]\mathrm{d}s\Big\|_{2}^{2}$
@@
    $\mathbb{E}_{\mathbb{P}}\Big[\Big\|\tilde{\gamma}_{t,u}-\hat{\gamma}_{t,u}\Big\|^{2}_{2}\Big]$
    $\leq C\,\mathbb{E}_{\mathbb{P}}\Big[\int_{0}^{t}\Big\|\tilde{\gamma}_{s,u}-\hat{\gamma}_{s,u}\Big\|_{2}^{2}\mathrm{d}s\Big]+\mathbb{E}_{\mathbb{P}}\Big[\Big\|\int_{0}^{t}\Big[\int_{0}^{1}(\tilde{w}(u,v)-\hat{w}(u,v))A^{\mathbb{E}_{\mathbb{P}}[\hat{\gamma}_{s,v}]}\mathrm{d}v,\tilde{\gamma}_{s,u}\Big]\mathrm{d}s\Big\|_{2}^{2}\Big],$
    $\leq C\Big(\int_{0}^{t}\mathbb{E}_{\mathbb{P}}\Big[\Big\|\tilde{\gamma}_{s,u}-\hat{\gamma}_{s,u}\Big\|_{2}^{2}\Big]\mathrm{d}s+\int_{0}^{t}\int_{0}^{1}\Big\|(\tilde{w}(u,v)-\hat{w}(u,v))A^{\mathbb{E}_{\mathbb{P}}[\hat{\gamma}_{s,v}]}\Big\|_{2}^{2}\mathrm{d}v\mathrm{d}s\Big),$
@@
    $\leq C\Big(\int_{0}^{t}\mathbb{E}_{\mathbb{P}}\Big[\Big\|\tilde{\gamma}_{s,u}-\hat{\gamma}_{s,u}\Big\|_{2}^{2}\Big]\mathrm{d}s+T\|a\|_{2}^{2}\int_{0}^{1}\Big|(\tilde{w}(u,v)-\hat{w}(u,v))\Big|^{2}\mathrm{d}v\Big).$
@@
    $\mathbb{E}_{\mathbb{P}}\Big[\int_{0}^{1}\Big\|\tilde{\gamma}_{t,u}-\hat{\gamma}_{t,u}\Big\|^{2}_{2}\mathrm{d}u\Big]\leq$ $C\Big(\int_{0}^{t}\mathbb{E}_{\mathbb{P}}\Big[\int_{0}^{1}\Big\|\tilde{\gamma}_{s,u}-\hat{\gamma}_{s,u}\Big\|_{2}^{2}\mathrm{d}u\Big]\mathrm{d}s$
@@
Remark 4.5 (Strong Solution).
In the case of perfect measurement where detector efficiency $\eta=1$, homogeneous graphon $w\equiv 1$ and bounded coupling operator $L$, the well-posedness in the strong sense was proved by a different method in [48]. An extension to the unbounded coupling operator was recently provided in [49]. For others results concerning well-posedness of stochastic Schrödinger equation, see [59, 35, 40] in infinite dimensional setting and [63] for the density operator equation in finite dimensional setting.
@@ 5 Propagation of chaos: Block-wise case
For clarity, we slightly re-adjust the framework and notation. Specifically, we consider $c$-packets of $N$ quantum particles interacting through a graph $G_{cN}=(V_{cN},E_{cN})$. The case where the blocks are not of equal size can be treated similarly. Indeed, we can consider $c$ packets, each containing $N_{i}$ particles, for $i\in[c]$, such that $\sum_{i=1}^{c}N_{i}=N,$ and $\lim_{N\to\infty}\frac{N_{i}}{N}=O(1)$. For simplicity of notation and exposition, we focus here on the case of equally-sized blocks, i.e., $N_{i}=N$ for all $i$.
@@
$\mathrm{d}\boldsymbol{\rho}^{cN}_{t}$ $=-\mathrm{i}\Big[\sum_{i_{l}=1_{1}}^{c_{N}}\mathbf{\tilde{H}}_{i_{l}}\!+\!\tfrac{1}{2cN}\sum_{i_{l}=1_{1}}^{c_{N}}\sum_{j_{l^{\prime}}=1_{1}}^{c_{N}}\!\xi_{j_{l^{\prime}}i_{l}}\mathbf{A}_{j_{l^{\prime}}i_{l}},\boldsymbol{\rho}^{cN}_{t}\Big]\!\mathrm{d}t$
$+\sum_{i_{l}=1_{1}}^{c_{N}}\big(\mathbf{L}_{i_{l}}\boldsymbol{\rho}^{cN}_{t}\mathbf{L}_{i_{l}}^{\dagger}-\tfrac{1}{2}\{\mathbf{L}_{i_{l}}^{\dagger}\mathbf{L}_{i_{l}},\boldsymbol{\rho}^{cN}_{t}\}\big)\mathrm{d}t$
$\qquad+\sum_{i_{l}=1_{1}}^{c_{N}}\Big(\mathbf{L}_{i_{l}}\boldsymbol{\rho}^{cN}_{t}+\boldsymbol{\rho}^{cN}_{t}\mathbf{L}_{i_{l}}^{\dagger}-\mathrm{t}\mathrm{r}\big((\mathbf{L}_{i_{l}}+\mathbf{L}_{i_{l}}^{\dagger})\boldsymbol{\rho}^{cN}_{t}\big)\boldsymbol{\rho}^{cN}_{t}\Big)\mathrm{d}W_{t}^{i_{l}}.$ (11)
@@
$\mathbf{O}_{i_{l}}$ $:=\overbrace{\mathbf{1}\otimes\dots\otimes\mathbf{1}}^{(i-1)N\text{ times}}\otimes\overbrace{\mathbf{1}\otimes\dots\otimes\underbrace{O}_{l-\text{th}}\otimes\dots\otimes\mathbf{1}}^{N\text{ times}}\otimes\overbrace{\mathbf{1}\otimes\dots\otimes\mathbf{1}}^{(c-i)N\text{ times}},$
@@
The interaction strength coefficient $\xi_{j_{l^{\prime}}i_{l}}:=\xi_{j_{l^{\prime}}i_{l}}^{cN}$ is derived from a step kernel function ${w}^{G_{cN}}$, and is constructed in the two following ways:
@@
$w(u,v)$ $:=\begin{cases}w_{ji},\quad(u,v)\in I_{j}\times I_{i},\\ 0,\quad\quad\,\,\,\textrm{otherwise.}\end{cases}$
@@
If the initial state is pure, i.e., $|\boldsymbol{\rho}^{cN}_{0}=\ket{\boldsymbol{\Psi}_{0}^{cN}}\bra{\boldsymbol{\Psi}_{0}^{cN}}$, then, by Proposition 2.2, the state remains pure i.e $|\boldsymbol{\rho}^{cN}_{t}=\ket{\boldsymbol{\Psi}_{t}^{cN}}\bra{\boldsymbol{\Psi}_{t}^{cN}}$ and can be written in the following pure state representation:
@@
Proposition 5.1 (Class‐wise symmetry).
@@
$\textrm{Law}\Big(\boldsymbol{\Psi}_{0}^{cN}\bigl(x_{1_{1}},\dots,x_{c_{N}}\bigr)\Big)$ $=\textrm{Law}\Big({\Psi}_{0}^{cN}\Bigl(x_{1_{\sigma^{1}(1)}},\dots,x_{1_{\sigma^{1}(N)}},\dots,x_{c_{\sigma^{c}(1)}},\dots,x_{c_{\sigma^{c}(N)}}\Bigr)\Big),$ (12)
for all permutations $\sigma^{1},\dots,\sigma^{c}\in\mathfrak{S}_{N}$. Then this symmetry is preserved in distribution by the $cN$-body Belavkin Equation (11) if the interaction coefficients $\{\xi_{j_{l^{\prime}}i_{l}}\}$ are invariants in law:
@@
Under Assumption 1 and Assumption 2, Equation (11) describes the dynamics of $c$-class of Bosons.
@@ 5.1 Limiting system
$\mathrm{d}\gamma_{t,u}$ $=-\mathrm{i}\Big[\tilde{H}+\int_{0}^{1}w(u,v)A^{\mathbb{E}_{\mathbb{P}}[\gamma_{t,v}]}\mathrm{d}v,\gamma_{t,u}\Big]\mathrm{d}t+\Big(L\gamma_{t,u}L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\gamma_{t,u}\}\Big)\mathrm{d}t$
@@
$\int_{0}^{1}{w}\Big(\frac{j-1}{c},v\Big)A_{j}^{\mathbb{E}_{\mathbb{P}}[{\gamma_{t,v}}]}\mathrm{d}v$ $=\sum_{i=1}^{c}\int_{0}^{1}\mathbb{I}_{\Big\{\frac{i-1}{c}\leq v<\frac{i}{c}\Big\}}{w}\Big(\frac{j-1}{c},v\Big)A_{j}^{\mathbb{E}_{\mathbb{P}}[{\gamma_{t,v}}]}\mathrm{d}v,$
$=\sum_{i=1}^{c}\int_{0}^{1}\mathbb{I}_{\Big\{\frac{i-1}{c}\leq v<\frac{i}{c}\Big\}}{w}\Big(\frac{j-1}{c},\frac{i-1}{c}\Big)A_{j}^{\mathbb{E}_{\mathbb{P}}[{\gamma_{t,v}}]}\mathrm{d}v,$
$=\sum_{i=1}^{c}\int_{\frac{i-1}{c}}^{\frac{i}{c}}{w}\Big(\frac{j-1}{c},\frac{i-1}{c}\Big)A_{j}^{\mathbb{E}_{\mathbb{P}}[{\gamma_{t,i}}]}\mathrm{d}v,$
$=\frac{1}{c}\sum_{i=1}^{c}{w}\Big(\frac{j-1}{c},\frac{i-1}{c}\Big)A_{j}^{\mathbb{E}_{\mathbb{P}}[{\gamma_{t,i}}]},$
$=\frac{1}{c}\sum_{i=1}^{c}w_{ji}A_{j}^{\mathbb{E}_{\mathbb{P}}[{\gamma_{t,i}}]}.$
@@
$\mathrm{d}\gamma_{t,j}$ $=-\mathrm{i}\Big[\tilde{H}+\frac{1}{c}\sum_{i=1}^{c}w_{ji}A^{\mathbb{E}_{\mathbb{P}}[\gamma_{t,i}]},\gamma_{t,j}\Big]\mathrm{d}t+\big(L\gamma_{t,j}L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\gamma_{t,j}\}\big)\mathrm{d}t$
$\qquad\qquad+\Big(L\gamma_{t,j}+\gamma_{t,j}L^{\dagger}-\mathrm{t}\mathrm{r}\big((L+L^{\dagger})\gamma_{t,j})\gamma_{t,j}\Big)\mathrm{d}W_{t}^{j},\;\;j\in[c].$ (13)